Tedder:

The area of the scale model of a garden is 15 square feet. The scale model is enlarged by a scale factor of 3 to create the actual garden. Which expression finds the area of the actual garden?

jakfishman:

still want help??

jakfishman:

and welcome to question cove

jhonyy9:

so suppose this garden has a rectangle form and you know that area of a rectangle you calcule using formula A = l * w so area rectangle equal the length time the width of the rectangle let be the scale model with dimensions of length an l and the width an w in this way the area you get A = l*w what given that A = 15 square feet and given again that to create the actual garden the scale model is enlarged by a scale factor of 3 so this mean that the length of the garden will be 3l and the width will be 3w and now using these new measures details of actual garden how you calcule area of actual garden using this above wrote formula ? hope helped understandably easy